A 2.8-kg disk is constrained horizontally but is free to move vertically. The disk is struck from below by a vertical jet of water. The speed and diameter of the water jet are 12 m/s and 25 mm at the nozzle exit. (a) Obtain a general expression for the speed of the water jet as a function of height, h. (b) Find the height to which the disk will rise and remain stationary.

M = 2.8 kg d=25 mm A Vo = 12 m/s
Answer #1

@ Bernoullin equation point z = Coustand onessure of both i jane + + n=2-31 wartagh v=rv5-agh speed of water jal calculat equ